# Representing rational integers by generalized quadratic forms over quadratic fields

### 11 Mar 2024 v1

@@ Abstract.
We acknowledge support by Czech Science Foundation (GAČR) grant 21-00420M and the Charles University Research Centre program UNCE/24/SCI/022.
@@ 1. Introduction
A positive definite quadratic form over $\mathbb{Z}$ is called universal if it represents every positive integer $n$. The history of universal quadratic forms begins with Lagrange’s proof of the Four Square Theorem in 1770: Every positive integer $n$ is of the form $x^{2}+y^{2}+z^{2}+w^{2}$. Diagonal quaternary universal forms were classified by Ramanujan [Ra], who stated that there exist $55$ such forms. Dickson [Di] corrected Ramanujan’s statement by pointing out that the form $x^{2}+2y^{2}+5z^{2}+5u^{2}$ does not represent $15$. Thus there exist exactly $54$ diagonal universal forms. Dickson [Di2] also proved various results on representation by non-diagonal quaternary forms.
The 90’s saw a proof of the celebrated Conway-Schneeberger 15-Theorem. A quadratic form having an integer matrix is called classical.
@@
[…]
The proof was simplified by Bhargava [Bh]. As a corollary, one obtains all classical quaternary forms – there are exactly $207$ of them. Bhargava and Hanke [BH] extended the result to non-classical forms by proving the 290-Theorem: If a positive definite quadratic form $Q$ over $\mathbb{Z}$ represents every positive integer $n\leq 290$, then $Q$ is universal. In fact, for $Q$ to be universal, it is enough that it represents every integer from a list of $29$ numbers, the largest number on the list being $290$. A corollary is that there are exactly $6436$ universal quaternary forms [BH, Theorem 4].
It is natural to extend the notion of universality from $\mathbb{Z}$ to the ring of integers $\mathcal{O}_{K}$ in a number field. Assume that the number field $K$ is totally real. A positive integral quadratic form over $K$ is called universal if it represents all totally positive integers in $K$. A version of the local-global principal over number fields proved by Hsia, Kitaoka, and Kneser [HKK] implies that a universal form over $K$ always exists. Maaß [Ma] showed that the sum of three squares $x^{2}+y^{2}+z^{2}$ is universal over $\mathbb{Q}(\sqrt{5})$. By a result of Siegel [Si], if a sum of squares is universal over $K$, then $K=\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$. In 1993, Chan, Kim, and Raghavan [CKR] proved that a ternary universal quadratic form exists over a real quadratic field $K=\mathbb{Q}(\sqrt{D})$ if and only if $D$ equals $2$, $3$, or $5$. Moreover, they identified all ternary universal forms over these fields up to equivalence. For forms of higher ranks, B. M. Kim [Ki] proved that there exist only finitely many $D$’s such that $\mathbb{Q}(\sqrt{D})$ admits a universal form in $7$ variables (see also [KKP2] for an explicit estimate on the size of $D$). On the other hand, a $8$-ary universal form exists over each $\mathbb{Q}(\sqrt{n^{2}-1})$ (when $n^{2}-1$ is squarefree) [Ki2]. Blomer and Kala [BK] showed: For any positive integer $r$, there are infinitely many quadratic fields $\mathbb{Q}(\sqrt{D})$ that do not have a universal lattice of rank $\leq r$. Lower bounds for the rank of a universal form depending on $D$ were obtained by Kala and Tinková [KT].
@@
In the same vein, one can study Hermitian forms over imaginary quadratic fields. We call a positive definite Hermitian form over an imaginary quadratic field $\mathbb{Z}$-universal if it represents every $n\in\mathbb{Z}_{\geq 1}$. Earnest and Khosravani [EK] proved that universal binary Hermitian forms exist over $K$ for only finitely many $m$, and identified all such forms in the case when $K$ has class number $1$. B. M. Kim, J. Y. Kim, and P.-S. Park [KKP] proved an analogy of the 15-Theorem for Hermitian quadratic forms: If a positive definite Hermitian form $H$ represents $1$, $2$, $3$, $5$, $6$, $7$, $10$, $13$, $14$, and $15$, then $H$ is universal. They also completely determined the minimal rank of a positive Hermitian form over all imaginary quadratic fields.
The object of study of our paper are generalized quadratic forms introduced recently by Browning, Pierce, and Schindler [BPS, Definition 1.1]. They can be defined for an arbitrary Galois extension $K$ of $\mathbb{Q}$ but we will consider them only over quadratic fields. A quadratic field $K=\mathbb{Q}(\sqrt{D})$ has two $\mathbb{Q}$-automorphisms: the identity and the conjugation map
$\tau:K\hookrightarrow K,a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
@@
$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in\mathbb{Z},\quad\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K},$
@@
We will prove these as Theorems 5.10, 5.11 and 5.12 in Section 5. In contrast with the situation for binary and ternary forms, it is not very difficult to construct a generalized form in four variables which is $\mathbb{Z}$-universal over every real quadratic field (and we do this later in Proposition 5.13). This provides a fairly complete picture of the minimal rank required for a generalized form to be $\mathbb{Z}$-universal.
@@
The rest of the paper is organized as follows. In Section 2, we collect the preliminaries about quadratic forms and quadratic fields. In Section 3, we find a minimal set of generators of the ring of $\mathbb{Q}$-valued generalized polynomials. The theory of generalized quadratic forms is developed in Section 4. We prove Theorems 1.2, 1.3 and 1.4 at the end of Section 5 and Theorem 6.6 in Section 6. We find all quaternary forms associated to $\mathbb{Z}$-universal binary generalized forms for $D\in\{2,3,6,7,10\}$ (and to classical ones for $D=5$) in Section 7.
@@ Acknowledgments
This research project was accomplished at the Student Number Theory Seminar at Charles University. We thank Vítězslav Kala for organizing the seminar. We also thank Nicolas Daans for suggesting the proof of Proposition 6.7 and Pavlo Yatsyna for his recommendation of the reference [Ca].
2. Preliminaries
Let $K$ be a number field and let $\mathcal{O}_{K}$ denote the ring of integers of $K$. Consider a polynomial $p$ in $n$ variables over $K$, i.e.,
$p(x_{1},x_{2},\dots,x_{n})=\sum_{\begin{subarray}{c}i_{1},\dots,i_{n}\geq 0\\ i_{1}+\cdots+i_{n}\leq k\end{subarray}}a_{i_{1},i_{2},\dots,i_{n}}x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$
where $k\in\mathbb{Z}_{\geq 0}$ and $a_{i_{1},i_{2},\dots,i_{n}}\in K$ for $i_{1},i_{2},\dots,i_{n}\in\{0,1,\dots,k\}$. We will make use of the multi-index notation. If $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})\in\mathbb{Z}_{\geq 0}^{n}$ is a multi-index, then we let $|\mathbf{i}|=\sum_{r=1}^{n}i_{r}$ be the sum of its components, and if $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$, then we let $\mathbf{x}^{\mathbf{i}}=x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$. The polynomial $p$ can be succinctly written as
$p(\mathbf{x})=\sum_{|\mathbf{i}|\leq k}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}.$
The lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$ is defined as follows: Let $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})$ and $\mathbf{j}=(j_{1},j_{2},\dots,j_{n})$. Then $\mathbf{i}<\mathbf{j}$ if and only if there exists $s\in\{1,2,\dots,n\}$ such that $i_{r}=j_{r}$ for $1\leq r<s$ and $i_{s}<j_{s}$.
A quadratic form $Q$ in $n$ variables over $K$ is a homogeneous polynomial of degree $2$, i.e.,
$Q(x_{1},x_{2},\dots,x_{n})=\sum_{1\leq i\leq j\leq n}a_{ij}x_{i}x_{j}$
where $a_{ij}\in K$. The matrix $M_{Q}$ of the form $Q$ is defined as
$M_{Q}=\begin{pmatrix}a_{11}&\frac{a_{12}}{2}&\dots&\frac{a_{1n}}{2}\\ \frac{a_{12}}{2}&a_{22}&\dots&\frac{a_{2n}}{2}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{a_{1n}}{2}&\frac{a_{2n}}{2}&\dots&a_{nn}\end{pmatrix}.$
The rank of $Q$ is the rank of the matrix $M_{Q}$ over $K$. If the coefficients $a_{ij}$ belong to $\mathcal{O}_{K}$, then $Q$ is called integral. If the matrix $M_{Q}$ has entries in $\mathcal{O}_{K}$, i.e., $a_{ij}\in 2\mathcal{O}_{K}$ for $i\neq j$, then $Q$ is called classical.
Let $Q$ be a quadratic form over $\mathbb{Q}$. We say that $Q$ represents $b\in\mathbb{Z}$ over $\mathbb{Z}$ if there exists $\mathbf{a}\in\mathbb{Z}^{n}$ such that $Q(\mathbf{a})=b$. In particular, according to our definition, every quadratic form represents 00. If there exists $\mathbf{a}\in\mathbb{Z}^{n}\setminus\{(0,0,\dots,0)\}$ such that $Q(\mathbf{a})=0$, then we say that $Q$ represents 00 non-trivially. The form $Q$ is called positive definite if $M_{Q}$ is a positive definite matrix and $Q$ is called universal if it is a positive definite integral form such that it represents every $b\in\mathbb{Z}_{\geq 1}$.
In the rest of the paper, we consider quadratic forms over $K$ only when $K=\mathbb{Q}$ or $K$ is a quadratic field. Let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. If we let
$\omega_{D}=\begin{cases}\sqrt{D},&\text{if }D\equiv 2,3\pmod{4},\\ \frac{1+\sqrt{D}}{2},&\text{if }D\equiv 1\pmod{4},\end{cases}$
then $(1,\omega_{D})$ is a basis of $\mathcal{O}_{K}$. The quadratic field $K$ has two embeddings into $\mathbb{C}$: the identity and the conjugation map
$\tau:K\hookrightarrow\mathbb{C},a+b\sqrt{D}\mapsto a-b\sqrt{D}.$
For $\alpha\in K$, we let $\operatorname{Tr}(\alpha)=\alpha+\tau(\alpha)$ and $\operatorname{N}(\alpha)=\alpha\tau(\alpha)$ be the trace and norm of $\alpha$, respectively. More generally, if $K$ is a number field with $r$ real and $2s$ complex embeddings
$\sigma_{1},\sigma_{2},\dots,\sigma_{r}:K\hookrightarrow\mathbb{R},\sigma_{r+1},\bar{\sigma}_{r+1},\dots,\sigma_{r+s},\bar{\sigma}_{r+s}:K\hookrightarrow\mathbb{C},$
then we let $V=\mathbb{R}^{r}\times\mathbb{C}^{s}$ and call $V$ the Minkowski space. The field $K$ embeds into $V$ via the map
$\sigma:K\hookrightarrow V,\alpha\mapsto(\sigma_{1}(\alpha),\dots,\sigma_{r+s}(\alpha)).$
We have $K\otimes_{\mathbb{Q}}\mathbb{R}\simeq V$ (see [Mi, p. 70]). In particular, $\sigma(K)$ is dense in $V$.
@@ 3. Generalized polynomials with rational values
where $k\in\mathbb{Z}_{\geq 0}$ and $\alpha_{i_{1},j_{1},\dots,i_{n},j_{n}}\in K$ for every $i_{1},j_{1},\dots,i_{n},j_{n}\in\{0,1,\dots,k\}$.
If $\mathbf{z}=(z_{1},z_{2},\dots,z_{n})$, then we let $\tau(\mathbf{z})=(\tau(z_{1}),\tau(z_{2}),\dots,\tau(z_{n}))$. Using the multi-index notation, we can write $g$ as
@@
First, suppose that $K$ is real ($r=2$ and $s=0$). We know that $\sigma(K)=\{(\alpha,\tau(\alpha)):\;\alpha\in K\}$ is dense in $V=\mathbb{R}^{2}$. Consequently, $\sigma(K)^{n}$ is dense in $\mathbb{R}^{2n}$. Because the polynomial $\tilde{g}$ vanishes on $\sigma(K)^{n}$, it vanishes on the whole set $\mathbb{R}^{2n}$, and hence $\tilde{g}$ is the zero polynomial.
Secondly, suppose that $K$ is imaginary ($r=0$ and $s=1$). In this case, $\sigma(K)=K$ is dense in $V=\mathbb{C}$. Thus $K^{n}$ is dense in $\mathbb{C}^{n}$ and the polynomial $\tilde{g}$ is zero on the set
@@
Let $g\in\Gamma_{K}^{1}$. We prove $g\in S_{1}$ by induction on $d=\deg(g)$. If $d=0$, then $g(z)=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ are contained in $S_{1}$. We write $g$ as
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$. By induction assumption, $h_{1}\in S_{1}$, and because we also know that $z\tau(z)\in S_{1}$, we get $g_{1}\in S_{1}$.
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$. By induction assumption, $h_{2}\in S_{2}$. We show $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$. If we let $\alpha_{d,0}=a+b\sqrt{D}$ for $a,b\in\mathbb{Q}$, then
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{1}\in S_{1}$. Because $(z+\tau(z))^{d}$ and $z\tau(z)$ also belong to $S_{1}$, we get $z^{d}+\tau(z)^{d}\in S_{1}$.
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree $<d$, hence $r_{2}\in S_{1}$. Because $\sqrt{D}(z-\tau(z))\in S_{1}$, we get $\sqrt{D}(z^{d}-\tau(z)^{d})\in S_{1}$. This proves $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$, thus $g_{2}\in S_{1}$.
@@
We proceed by induction on $n$. For $n=1$, this is Lemma 3.5. Suppose that $n\geq 2$ and the statement is true for polynomials with $<n$ variables.
Let $g\in\Gamma_{K}^{n}$. We show $g\in S_{n}$ by induction on $d=\deg(g)$. If $d=0$, then $g(\mathbf{z})=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree $<d$ in $n$ variables are contained in $S_{n}$. The polynomial $g$ is a $\mathbb{Q}$-linear combination of the elements in the basis from Lemma 3.6. Thus it is enough to show that the basis elements of degree $d$ belong to $S_{n}$.
@@
If $i_{n}=0$ and $j_{n}=0$, then the variable $z_{n}$ does not appear in $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$, hence $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ by induction assumption.
@@
Because $h(\mathbf{z})$ has degree $d-1$, we can apply induction on $h(\mathbf{z})+\tau(h(\mathbf{z}))$ and $\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))$. Since
@@
If $i_{n}=0$, then $\mathbf{g}_{ii}$ is a polynomial in $n-1$ variables, hence $\mathbf{g}_{ii}\in S_{n}$ by induction assumption.
@@
The generalized polynomial $h$ has degree $d-2$, hence $h\in S_{n}$ by induction assumption, and $g_{\mathbf{i}\mathbf{i}}\in S_{n}$ follows. ∎

### 21 Mar 2026 v2

@@ Abstract.
We acknowledge support by Czech Science Foundation (GAČR) grant 26-20514S, and Charles University programmes PRIMUS/24/SCI/010 and UNCE/24/SCI/022, and GAUK projects 134824 and 236225.
@@ 1. Introduction
A positive definite quadratic form over $\mathbb{Z}$ is called universal if it represents every positive integer $n$. The history of universal quadratic forms begins with Lagrange’s proof of the Four Square Theorem in 1770: Every positive integer $n$ is of the form $x^{2}+y^{2}+z^{2}+w^{2}$. Diagonal quaternary universal forms were classified by Ramanujan [Ra], who stated that there exist $55$ such forms. Dickson [Di1] corrected Ramanujan’s statement by pointing out that the form $x^{2}+2y^{2}+5z^{2}+5u^{2}$ does not represent $15$. Thus there exist exactly $54$ diagonal universal forms. Dickson [Di2] also proved various results on representation by non-diagonal quaternary forms.
In 1993, John H. Conway and W. A. Schneeberger proved the celebrated 15-Theorem. A quadratic form having an integer matrix is called classical.
@@
$1,2,3,5,6,7,10,14,\text{and }15.$
The proof was simplified by Bhargava [Bh]. As a corollary, one obtains all universal classical quaternary forms – there are exactly $207$ of them. Bhargava and Hanke [BH] extended the result to non-classical forms by proving the 290-Theorem: If a positive definite quadratic form $Q$ over $\mathbb{Z}$ represents every positive integer $n\leq 290$, then $Q$ is universal. In fact, for $Q$ to be universal, it is enough that it represents every integer from a list of $29$ numbers, the largest number on the list being $290$. A corollary is that there are exactly $6436$ universal quaternary forms [BH, Theorem 4].
It is natural to extend the notion of universality from $\mathbb{Z}$ to the ring of integers $\mathcal{O}_{K}$ in a number field. Assume that the number field $K$ is totally real. A positive integral quadratic form over $K$ is called universal if it represents all totally positive integers in $K$. A version of the local-global principal over number fields proved by Hsia, Kitaoka, and Kneser [HKK] implies that a universal form over $K$ always exists. Maaß [Ma] showed that the sum of three squares $x^{2}+y^{2}+z^{2}$ is universal over $\mathbb{Q}(\sqrt{5})$. By a result of Siegel [Si], if a sum of squares is universal over $K$, then $K=\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$. In 1993, Chan, Kim, and Raghavan [CKR] proved that a ternary universal quadratic form exists over a real quadratic field $K=\mathbb{Q}(\sqrt{D})$ if and only if $D$ equals $2$, 33, or $5$. Moreover, they identified all ternary universal forms over these fields up to equivalence. For forms of higher ranks, B. M. Kim [Ki1] proved that there exist only finitely many $D$’s such that $\mathbb{Q}(\sqrt{D})$ admits a universal form in $7$ variables (see also [KKP2] for an explicit estimate on the size of $D$). On the other hand, a $8$-ary universal form exists over each $\mathbb{Q}(\sqrt{n^{2}-1})$ whenever $n^{2}-1$ is squarefree [Ki2]. Blomer and Kala [BK1] showed: For any positive integer $r$, there are infinitely many quadratic fields $\mathbb{Q}(\sqrt{D})$ that do not have a universal lattice of rank $\leq r$. Lower bounds for the rank of a universal form depending on $D$ were obtained by Kala and Tinková [KT].
@@
In the same vein, one can study Hermitian forms over imaginary quadratic fields. We call a positive definite Hermitian form over an imaginary quadratic field $\mathbb{Z}$-universal if it represents every $n\in\mathbb{Z}_{\geq 1}$. Earnest and Khosravani [EK] proved that universal binary Hermitian forms exist over $K=\mathbb{Q}(\sqrt{m})$ for only finitely many $m$, and identified all such forms in the case when $K$ has class number $1$. B. M. Kim, J. Y. Kim, and P.-S. Park [KKP1] proved an analogy of the 15-Theorem for Hermitian quadratic forms: If a positive definite Hermitian form $H$ represents $1$, 22, 33, 55, 66, 77, $10$, $13$, $14$, and $15$, then $H$ is universal. They also completely determined the minimal rank of a positive Hermitian form over all imaginary quadratic fields.
Our paper concerns generalized quadratic forms introduced recently by Browning, Pierce, and Schindler [BPS, Definition 1.1]. They can be defined for an arbitrary Galois extension $K$ of $\mathbb{Q}$ but we will consider them only over quadratic fields. A quadratic field $K=\mathbb{Q}(\sqrt{D})$ has two $\mathbb{Q}$-automorphisms: the identity and the conjugation map
$\tau:K$ $\hookrightarrow K$
$a+b\sqrt{D}$ $\mapsto a-b\sqrt{D}.$
@@
$G(\alpha_{1},\alpha_{2},\dots,\alpha_{n})\in\mathbb{Z},\quad\forall\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathcal{O}_{K},$
@@
We will prove these as Theorems 4.10, 4.11 and 4.12 in Section 4. In contrast with the situation for binary and ternary forms, it is not very difficult to construct a generalized form in four variables which is $\mathbb{Z}$-universal over every real quadratic field (and we do this later in Proposition 4.13). This provides a fairly complete picture of the minimal rank required for a generalized form to be $\mathbb{Z}$-universal.
@@
The rest of the paper is organized as follows. In Section 2, we find a minimal set of generators of the ring of $\mathbb{Q}$-valued generalized polynomials. The theory of generalized quadratic forms is developed in Section 3. We prove Theorems 1.2, 1.3 and 1.4 at the end of Section 4 and Theorem 5.1 in Section 5. In the appendix, we give the proofs of Lemmata 4.5-4.9 in Section A4 and a more detailed approach to the proof of Theorem 5.1 in Section A5. In Section A6 we explicitly show all quaternary forms associated to $\mathbb{Z}$-universal binary generalized forms for $D\in\{2,3,6,7,10\}$ (and to classical ones for $D=5$).
@@ Acknowledgments
This research project was accomplished at the Student Number Theory Seminar at Charles University. We thank Nicolas Daans, Vítězslav Kala and Pavlo Yatsyna for their helpful suggestions and the anonymous referee for their constructive feedback.
@@ 2. Generalized polynomials with rational values
Throughout the article let $K=\mathbb{Q}(\sqrt{D})$ where $D\in\mathbb{Z}\setminus\{0,1\}$ is squarefree. If we let
[…]
then $(1,\omega_{D})$ is a basis of $\mathcal{O}_{K}$. For $\alpha\in K$, we let $\operatorname{Tr}(\alpha)=\alpha+\tau(\alpha)$ and $\operatorname{N}(\alpha)=\alpha\tau(\alpha)$ be the trace and norm of $\alpha$, respectively.
@@
where $k\in\mathbb{Z}_{\geq 0}$, $\alpha_{i_{1},j_{1},\dots,i_{n},j_{n}}\in K$ for every $i_{1},j_{1},\dots,i_{n},j_{n}\in\{0,1,\dots,k\}$ and $\tau:K\hookrightarrow\mathbb{C}$ where $\tau(a+b\sqrt{D})=a-b\sqrt{D}$ is the conjugation map.
We will often use the multiindex notation, where $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})\in\mathbb{Z}_{\geq 0}^{n}$ is a multiindex, $|\mathbf{i}|=\sum_{r=1}^{n}i_{r}$ is the sum of its components, and for $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ we let $\mathbf{x}^{\mathbf{i}}=x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{n}^{i_{n}}$. The lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$ is defined as follows: Let $\mathbf{i},\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}$, $\mathbf{i}=(i_{1},i_{2},\dots,i_{n})$ and $\mathbf{j}=(j_{1},j_{2},\dots,j_{n})$. Then $\mathbf{i}<\mathbf{j}$ if and only if there exists $s\in\{1,2,\dots,n\}$ such that $i_{r}=j_{r}$ for $1\leq r<s$ and $i_{s}<j_{s}$.
Hence if $\mathbf{z}=(z_{1},z_{2},\dots,z_{n})$, then we let $\tau(\mathbf{z})=(\tau(z_{1}),\tau(z_{2}),\dots,\tau(z_{n}))$ and we can write the generalized polynomial $g$ as
@@
First, suppose that $K$ is real. We know that $\sigma(K)=\{(\alpha,\tau(\alpha)):\;\alpha\in K\}$ is dense in $V=\mathbb{R}^{2}$. Consequently, $\sigma(K)^{n}$ is dense in $\mathbb{R}^{2n}$. Because the polynomial $\tilde{g}$ vanishes on $\sigma(K)^{n}$, it vanishes on the whole set $\mathbb{R}^{2n}$, and hence $\tilde{g}$ is the zero polynomial.
Secondly, suppose that $K$ is imaginary. In this case, $\sigma(K)=K$ is dense in $V=\mathbb{C}$. Thus $K^{n}$ is dense in $\mathbb{C}^{n}$ and the polynomial $\tilde{g}$ is zero on the set
@@
Let $g\in\Gamma_{K}^{1}$. We prove $g\in S_{1}$ by induction on $d=\deg(g)$. If $d=0$, then $g(z)=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree <d<d are contained in $S_{1}$. We write $g$ as
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d. By the inductive hypothesis, $h_{1}\in S_{1}$, and because we also know that $z\tau(z)\in S_{1}$, we get $g_{1}\in S_{1}$.
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d. By the inductive hypothesis, $h_{2}\in S_{2}$. We show $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$. If we let $\alpha_{d,0}=a+b\sqrt{D}$ for $a,b\in\mathbb{Q}$, then
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d, hence $r_{1}\in S_{1}$. Because $(z+\tau(z))^{d}$ and $z\tau(z)$ also belong to $S_{1}$, we get $z^{d}+\tau(z)^{d}\in S_{1}$.
@@
is a $\mathbb{Q}$-valued generalized polynomial of degree <d<d, hence $r_{2}\in S_{1}$. Because $\sqrt{D}(z-\tau(z))\in S_{1}$, we get $\sqrt{D}(z^{d}-\tau(z)^{d})\in S_{1}$. This proves $\alpha_{d,0}z^{d}+\tau(\alpha_{d,0})\tau(z)^{d}\in S_{1}$, thus $g_{2}\in S_{1}$.
@@
We proceed by induction on $n$. For $n=1$, this is Lemma 2.5. Suppose that $n\geq 2$ and the statement is true for polynomials with <n<n variables.
Let $g\in\Gamma_{K}^{n}$. We show $g\in S_{n}$ by induction on $d=\deg(g)$. If $d=0$, then $g(\mathbf{z})=a\in\mathbb{Q}$ is a constant polynomial. Thus we assume that $d\geq 1$ and that the $\mathbb{Q}$-valued generalized polynomials of degree <d<d in $n$ variables are contained in $S_{n}$. The polynomial $g$ is a $\mathbb{Q}$-linear combination of the elements in the basis from Lemma 2.6. Thus it is enough to show that the basis elements of degree $d$ belong to $S_{n}$.
@@
If $i_{n}=0$ and $j_{n}=0$, then the variable $z_{n}$ does not appear in $g_{\mathbf{i}\mathbf{j}}$ and $g_{\mathbf{j}\mathbf{i}}$, hence $g_{\mathbf{i}\mathbf{j}},g_{\mathbf{j}\mathbf{i}}\in S_{n}$ by the inductive hypothesis.
@@
Because $h(\mathbf{z})$ has degree d1d-1, we can apply induction on $h(\mathbf{z})+\tau(h(\mathbf{z}))$ and $\sqrt{D}(h(\mathbf{z})-\tau(h(\mathbf{z})))$. Since
@@
If $i_{n}=0$, then $\mathbf{g}_{ii}$ is a polynomial in n1n-1 variables, hence $\mathbf{g}_{ii}\in S_{n}$ by the inductive hypothesis.
@@
The generalized polynomial $h$ has degree d2d-2, hence $h\in S_{n}$ by the inductive hypothesis, and $g_{\mathbf{i}\mathbf{i}}\in S_{n}$ follows. ∎